OpenStudy (moyabote):

use substitution to solve 3X+Y=6 4X+2Y=8

pooja195 (pooja195):

Hi there! Welcome To OpenStudy This question belongs in the math section however i'll help! ^_^ Ok any ideas on where to start? :-)

OpenStudy (moyabote):

Hi Thank you, well no really...

pooja195 (pooja195):

No worries :-) Ok, so are you fimiliar with the substitution method?

OpenStudy (moyabote):

I'm trying to understanding it

pooja195 (pooja195):

Alright, 3X+Y=6 Do you know how to isolate y?

OpenStudy (moyabote):

I will do Y=3X+6?

pooja195 (pooja195):

almost! y=-3x+6 Remember the we subtracted 3x from both sides... Now we have: y=-3x+6 <---equation 1 4X+2Y=8 <---equation 2 We know the value of y so we can plug equation 1 into equation 2 4x+2(-3x+6)=8 try solving for x

OpenStudy (moyabote):

Thank you, may sense 4X-6X+6=8 -2X=2 X=1

OpenStudy (moyabote):

Is that correct?

pooja195 (pooja195):

Not right... 4x+2(-3x+6)=8 4x-6x+12=8 Remember that you need to distribute the 2 4x-6x+12=8 solve for x from here

OpenStudy (moyabote):

-2X+12=8 -2X=-4 X=-4 ?

OpenStudy (usukidoll):

A negative number divided by a negative number becomes positive

pooja195 (pooja195):

-2x+12=8 subtract 12 from both sides

pooja195 (pooja195):

@UsukiDoll reguardless the answer is wrong.

OpenStudy (moyabote):

:(

OpenStudy (usukidoll):

At least the user tried...geez

OpenStudy (moyabote):

Ok let me try again

pooja195 (pooja195):

Keep trying! :) And @UsukiDoll I think the user would rather get proper help, would you like me to say the answer is correct when it isn't? Anyways @moyabote -2x+12=8 subtract 12 from both sides

OpenStudy (usukidoll):

And ugh 8-12 is -4 which I saw earlier 4x -6x is -2x which I saw earlier as well it was almost right but the division part was wrong.

OpenStudy (moyabote):

X=2 ?

pooja195 (pooja195):

Yep perfect :)

pooja195 (pooja195):

Now we take x=2 and plug it into equation 1 :) y=-3(2)+6 now just solve for y :) and we're done B)

OpenStudy (moyabote):

Thank you so much, by the way, I'm a concern mother trying to understand so I can explain to my child, but I'm from Colombia and I finished high school in 1990, and the method is very different. Thank you for your help, I really appreciate it :)

pooja195 (pooja195):

No problem hope you continue to use the site :D And you can also help your child by using this https://www.mathsisfun.com/algebra/substitution.html its more visual

OpenStudy (moyabote):

thank you so much

pooja195 (pooja195):

yw ^_^